MERRIC IS BUYING A BOARD TO USE AS A DIAGONAL BRACE ON A RECTANGULAR GATE. THE GATE IS 36 INCHES AND 42 INCHES TALL. WHAT LENGTH OF BOARD SHOULD BE USED FOR THE BRACE?
Answer by checkley75(3666)

36^2=42^2=b^2
1296+1764=b^2
3060=b^2
55.317266=b or the hypotenuse aka. the brace.